# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ Abstract
We implement the geometric method proposed in ([7][3][12]) to analytically predict the sequence of bifurcations leading to a change of stability and/or the appearance of new periodic orbits in the secular 3D planetary three body problem. Stemming from the analysis in [13], we examine various normal form models as regards the extent to which they lead to a phase space dynamics qualitatively similar as that in the complete system. For fixed total angular momentum, the phase space in Hopf variables is the 3D sphere, and the complete sequence of bifurcations of new periodic orbits can be recovered through formulas yielding the tangencies or degenerate intersections between the sphere and the surfaces of a constant second integral of motion of the normal form flow. In particular, we find the critical values of the second integral giving rise to pitchfork and saddle-node bifurcations of new periodic orbits in the system. This analysis renders possible to predict the most important structural changes in the phase space, as well as the emergence of new possible stable periodic planetary orbital configurations which can take place as the mutual inclination between the two planets is allowed to increase.
@@ 1 Introduction
The historical importance of the planetary three body problem in the conceptual growth of dynamical systems’ theory in general cannot be overemphasized. Stemming from Laplace and Lagrange’s investigations on the ‘secular’ (long-term) evolution of the planetary orbits in our own solar system ( [8][9]), the model was shown by Poincaré to serve as an archetype for the development of key ideas in dynamical systems, including the importance of periodic orbits in sculpting the phase-space structure, as well as the concepts of (near-)integrability and chaos ( [17]).
In recent years, new interest in the problem has emerged owing to its close relevance to understanding planetary motions in extrasolar planetary systems. We refer the reader to the introduction in [13] for an extended literature review. Most recent studies stem from the planar case, i.e., where the orbits of all three bodies are co-planar. A basic known configuration of planetary orbits in such systems is the one of apsidal corotation resonance (ACR). In this configuration, the two planets move in osculating Keplerian ellipses whose pericenters undergo precession both with the same ‘secular’ (i.e., slow) frequency, remaining always either anti-aligned (hereafter ‘state A’) or always aligned (‘state B’). Mathematically, the ACR states are periodic orbits corresponding to fixed points of the secular Hamiltonian of the system (see, for example, [10][1][11]). This is an integrable Hamiltonian produced from the Hamiltonian (1) taking the restriction to co-planar orbits, and averaging with respect to the planetary mean anomalies (see section 2).
@@
$\mathrm{AMD}=L_{2}+L_{3}-L_{z}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (2)
In Eq. (2) $L_{2}$, $L_{3}$ are the angular momenta of the circular orbits of semi-major axes $a_{2},a_{3}$ equal to those of the two planets, and $L_{z}$ the modulus of the total angular momentum normal to the system’s Laplace plane. The lack of a third global integral implies that the system can in principle admit also chaotic solutions. As shown in the recent work [13], chaotic motions actually occupy a substantial volume in phase space as the mutual inclination of the two planets approaches the critical limit of birth of the so-called ‘Lidov-Kozai’ regime, where the configuration with one of the two planets in circular orbit changes stability character, turning from stable to unstable.
@@
$\mathcal{H}_{sec}=\mathcal{H}_{planar}(\mathbf{X},\mathbf{Y})+\mathcal{H}_{space}(\mathbf{X},\mathbf{Y};\mathrm{AMD})\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (3)
@@
$\mathcal{H}_{sec}=\mathcal{H}_{int}(\mathbf{X},\mathbf{Y})+\mathcal{H}_{1,space}(\mathbf{X},\mathbf{Y};\mathrm{AMD})\leavevmode\nobreak\ \leavevmode\nobreak\ ,$ (6)
@@
$\mathcal{H}_{int}^{(1)}=\mathcal{H}_{int}+{1\over 2}<\{\mathcal{H}_{1,space},\chi\}>\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (7)
[…]
@@
iii) Our third model considered, ~$\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$, is similar to the secular normal form $\mathcal{H}_{int}^{(1)}$, but obtained within the framework of the octupole approximation of the initial Hamiltonian truncated up to fourth order in the orbital eccentricities. Our interest in such a model, which becomes precise in cases of so-called ‘hierarchical’ systems (planets in nearly circular orbits with $a_{2}/a_{3}<<1$), is motivated by the simplicity in the form of the Hamiltonian, which allows for analytical formulas to be arrived at regarding the dependence of the sequence of bifurcations on the system’s parameters.
@@
Focusing, now, on the above three examples of integrable models approximating the 3D secular dynamics, we implement the geometric method exposed in detail in Section 3 to obtain theoretical predictions for the sequences of bifurcations of periodic orbits arising in each model. New periodic orbits appear stemming from either of the ACR modes A or B, through both saddle-node and pitchfork bifurcations. The sequences are parameterized altering the values of each of the integrals of the energy $\mathcal{E}$ and angular momentum $\sigma_{0}$. The method relies on the use of suitable Hopf variables $\sigma_{1},\sigma_{2},\sigma_{3}$ defined so that, for fixed $\sigma_{0}$, the corresponding reduced manifold in phase space is the 3-sphere $\sigma_{1}^{2}+\sigma_{2}^{2}+\sigma_{3}^{2}=\sigma_{0}^{2}$. All quasi-periodic motions of the system are represented by the curves of transverse intersections of the the 2D surfaces of constant energy (e.g. $\mathcal{H}_{int}(\sigma_{0},\sigma_{1},\sigma_{3})=\mathcal{E}$) with a sphere of fixed $\sigma_{0}$. All periodic orbits of the system can then be classified in terms of two different kinds of geometric connections between constant energy surfaces and the spheres of different radii $\sigma_{0}$: i) tangencies, and ii) transverse intersections, where the constant energy condition leads to a surface of degenerate form. Section 3 makes a systematic classification of these cases, the corresponding type of stability of the resulting periodic orbits, as well as the various possibilities for the resulting sequences of bifurcations, depending on the chosen integrable model. Whenever possible, we also provide the corresponding analytical formulas pertinent to the above analysis.
@@
To verify the validity of the formulas deduced in section 3, Section 4 implements the method to each of the models $\mathcal{H}_{int}$, $\mathcal{H}_{int}^{(1)}$, or ~$\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$. Besides showing the ability of the method to predict the correct sequence of bifurcations observed in each model, a key result is that we find that even small differences in the chosen model can lead to drastic differences as regards the observed sequence of bifurcations. We explain the geometric origin of such differences. Comparing to the initial non-integrable Hamiltonian (Eq. (3)), we find that the model $\mathcal{H}_{int}^{(1)}$, which represents a higher order normal form for the secular Hamiltonian, yields also the best predictions as regards both the correct thresholds in parameter values and the form of the sequence of bifurcations. Of course, the question of how precisely a secular normal form model can approximate the true dynamics can only be answered through case by case study, i.e, studying various cases of exoplanetary systems. While this is beyond our present scope, we stress that the method exposed in Section 3 is equally applicable to all such cases.
@@ 2.1 Secular Hamiltonian
Following, now, section $2.3$ of [13], we first Jacobi-reduce the Hamiltonian (8), introducing a ‘book-keeping symbol’ which keeps track of the order of smallness in eccentricity and mutual inclination of all the terms in the Hamiltonian expansion. This leads to the decomposition:
$\mathcal{H}_{sec}(w_{2},w_{3},W_{2},\,W_{3};L_{z})=\mathcal{H}_{planar}(w_{2}-w_{3},W_{2},W_{3})+\mathcal{H}_{space}(w_{2},w_{3},W_{2},W_{3};L_{z})\,,$ (10)
where
@@
are modified Delaunay variables, with $j=2,3$, and
@@
$H_{j}=G_{j}\cos(\mathrm{i}_{j})\,,$ $h_{j}=\Omega_{j}\,.$
The spatial part $\mathcal{H}_{space}$ can be split as $\mathcal{H}_{space}=\mathcal{H}_{0,{space}}+\mathcal{H}_{1,{space}}$, where $\mathcal{H}_{0,{space}}$ contains terms independent of the angles or depending trigonometrically on the difference $w_{2}-w_{3}=\omega_{3}-\omega_{2}$, and $\mathcal{H}_{1,{space}}$ contains terms depending trigonometrically on combinations of the angles $w_{2},w_{3}$ other than their difference. Then we write
@@
The first two terms in the above expression give rise to a 2 degrees of freedom integrable Hamiltonian
@@
(where the $\mathrm{AMD}$ is defined as in Eq. (2)), whose second integral is $W_{2}+W_{3}$.
@@ 2.2 Numerical phase portraits
[…]
@@
$X_{j}=-\sqrt{2\,W_{j}}\cos(w_{j})\,,$ $Y_{j}=\sqrt{2\,W_{j}}\sin(w_{j})\,,$ $j=2,3\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (16)
@@
$\cos(\mathrm{i}_{mut})$ $=\frac{L_{z}^{2}-\Lambda_{2}^{2}-\Lambda_{3}^{2}+\Lambda_{2}^{2}\,\mathrm{e}_{2}^{2}+\Lambda_{3}^{2}\,\mathrm{e}_{3}^{2}}{2\Lambda_{2}\Lambda_{3}\sqrt{1-\mathrm{e}_{2}^{2}}\sqrt{1-\mathrm{e}_{3}^{2}}}=\frac{L_{z}^{2}-G_{2}^{2}-G_{3}^{2}}{2G_{2}G_{3}}\leavevmode\nobreak\ \leavevmode\nobreak$ (17)
@@
$\mathrm{i}_{max}=\cos^{-1}\left(\frac{L_{z}^{2}-\Lambda_{2}^{2}-\Lambda_{3}^{2}}{2\Lambda_{2}\Lambda_{3}}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (18)
@@
Figure 1: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ in the plane $(X_{2},Y_{2})$ with $\mathrm{AMD}$ fixed and different values of energy. The surfaces of section are computed by a numerical integration of trajectories in the Hamiltonian truncated at multipolar degree $N_{P}=5$, order $N_{bk}=12$ in the eccentricities, and energies (from left to right) $\mathcal{E}=-6.77\cdot 10^{-5},-2.53\cdot 10^{-5},-1.92\cdot 10^{-5},-1.17\cdot 10^{-5}$,$-2.61\cdot 10^{-6}$.
@@
Figure 1 allows to recognize one out of all in-principle possible chains of bifurcations that can emerge in 3D secular models of the type discussed above. Such models may differ in the parameters (masses and semi-major axes of the planets, value of the AMD) as well as on the truncation orders (in the multipoles and/or the orbital eccentricities and inclinations). Our aim in the next section is to provide a systematic method to predict which out of all possibilities actually takes place in a system with given parameters. In the case of Fig. 1, a quick description of the numerically obtained sequence of bifurcations is as follows: for low $\mathcal{E}$, i.e., nearly planar orbits (left most panel in 1) the phase portrait exhibits two stable fixed points which correspond to the ACR modes A (right fixed point) and B (left fixed point, reversed with respect to the left-right position of the same fixed points in figure 9 of [13]). In between the centers A and B we have a continuous transition of quasi-periodic orbits represented in the surface of section by closed curves surrounding either the fixed point A or B. No unstable points or separatrices exist between the stable points, while the apparent separation of the two domains is only an artifact of the projection of the phase space, whose real topology is the one of the 2-sphere (see below), to the plane $(X_{2},Y_{2})$. The second panel in Fig. 1 shows, now, the birth of the families called ‘C1’ (unstable) and ‘C2’ (stable) in [13] through a saddle-node bifurcation. The unstable orbit C1 quickly becomes stable through a pitchfork bifurcation giving rise to two ‘off-axis’ ($Y_{2}\neq 0$) unstable orbits which later (as the energy $\mathcal{E}$ increases) join each other through an inverse pitchfork bifurcation turning the left fixed point of the middle panel of Fig. 1 from stable to unstable. For still higher energy $\mathcal{E}$, the C2 orbit finally turns also to unstable via the ‘Lidov-Kozai’ mechanism. Note that in this last regime a large volume of the phase space is dominated by chaotic motions, hence excluded by any of the integrable models used in the sequel to approximate the flow under the complete Hamiltonian.
Refer to captionRefer to caption
Figure 2: Top frame: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}}(\mathcal{E};\mathrm{AMD})$ for the integrable Hamiltonian model $\mathcal{H}_{int}$ at the values of the energies (from top to bottom, left to right) $\mathcal{E}=-6.77\cdot 10^{-5},-1.8\cdot 10^{-5},-1.74\cdot 10^{-5},-1.7\cdot 10^{-5},-1.67\cdot 10^{-5},-1.6\cdot 10^{-5}$,$-1.43\cdot 10^{-5},-4.05\cdot 10^{-7}$. Bottom frame: Phase portraits by contour plots $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0};\mathrm{AMD})$ with decreasing values of $\sigma_{0}$ (from top to bottom, left to right) $\sigma_{0}=1.62\cdot 10^{-2},6.15\cdot 10^{-3},5.93\cdot 10^{-3},5.85\cdot 10^{-3},5.78\cdot 10^{-3},5.62\cdot 10^{-3},5.304\cdot 10^{-3},1.74\cdot 10^{-4}$. The two alternative representations of the phase portraits, corresponding to $\mathcal{P_{Z}}(\mathcal{E})$ (top) or $\mathcal{P_{Z}}(\sigma_{0})$ (bottom), for $\mathcal{Z}=\mathcal{H}_{int}$, yield an equivalent bifurcation sequence.

### v3

@@ Abstract
We implement the geometric method proposed in ([9], [3], [16]) to analytically predict the sequence of bifurcations leading to a change of stability and/or the appearance of new periodic orbits in the secular 3D planetary three body problem. Stemming from the analysis in [17], we examine various normal form models as regards the extent to which they lead to a phase space dynamics qualitatively similar as that in the complete system. For fixed total angular momentum, the phase space in Hopf variables is the 3D sphere, and the complete sequence of bifurcations of new periodic orbits can be recovered through formulas yielding the tangencies or degenerate intersections between the sphere and the surfaces of a constant second integral of motion of the normal form flow. In particular, we find the critical values of the second integral giving rise to pitchfork and saddle-node bifurcations of new periodic orbits in the system. This analysis renders possible to predict the most important structural changes in the phase space, as well as the emergence of new possible stable periodic planetary orbital configurations which can take place as the mutual inclination between the two planets is allowed to increase.
@@ 1 Introduction
The historical importance of the planetary three body problem in the conceptual growth of dynamical systems’ theory in general cannot be overemphasized. Stemming from Laplace and Lagrange’s investigations on the ‘secular’ (long-term) evolution of the planetary orbits in our own solar system ([10], [11]), the model was shown by Poincaré to serve as an archetype for the development of key ideas in dynamical systems, including the importance of periodic orbits in sculpting the phase-space structure, as well as the concepts of (near-)integrability and chaos ([23]).
In recent years, new interest in the problem has emerged owing to its close relevance to understanding planetary motions in extrasolar planetary systems. We refer the reader to the introduction in [17] for an extended literature review. Most recent studies stem from the planar case, i.e., where the orbits of all three bodies are co-planar. A basic known configuration of planetary orbits in such systems is the one of apsidal corotation resonance (ACR). In this configuration, the two planets move in osculating Keplerian ellipses whose pericenters undergo precession both with the same ‘secular’ (i.e., slow) frequency, remaining always either anti-aligned (hereafter ‘state A’) or always aligned (‘state B’). Mathematically, the ACR states are periodic orbits corresponding to fixed points of the secular Hamiltonian of the system (see, for example, [12], [1], [13]). This is an integrable Hamiltonian produced from the Hamiltonian (1) taking the restriction to co-planar orbits, and averaging with respect to the planetary mean anomalies (see section 2).
@@
$\mathrm{AMD}=L_{2}+L_{3}-L_{z}\penalty 10000\ \penalty 10000\ .$ (2)
In Eq. (2) $L_{2}$, $L_{3}$ are the angular momenta of the circular orbits of semi-major axes $a_{2},a_{3}$ equal to those of the two planets, and $L_{z}$ the modulus of the total angular momentum normal to the system’s Laplace plane. Ignoring the precession of the common nodal line of the planetary orbits in the Laplace plane, the conservation of the angular momentum essentially means that the system can be formally treated as with 2 DOF. The lack of a third global integral implies that the system can admit also chaotic solutions. As shown in the recent work [17], chaotic motions actually occupy a substantial volume in phase space as the mutual inclination of the two planets approaches the critical limit of birth of the so-called ‘Lidov-Kozai’ regime, where the configuration with one of the two planets in circular orbit changes stability character, turning from stable to unstable.
@@
$\mathcal{H}_{sec}=\mathcal{H}_{planar}(\mathbf{X},\mathbf{Y})+\mathcal{H}_{space}(\mathbf{X},\mathbf{Y};\mathrm{AMD})\penalty 10000\ \penalty 10000\ .$ (3)
@@
$\mathcal{H}_{sec}=\mathcal{H}_{int}(\mathbf{X},\mathbf{Y})+\mathcal{H}_{1,space}(\mathbf{X},\mathbf{Y};\mathrm{AMD})\penalty 10000\ \penalty 10000\ ,$ (6)
@@
$\mathcal{H}_{int}^{(1)}=\mathcal{H}_{int}+{1\over 2}<\{\mathcal{H}_{1,space},\chi\}>\penalty 10000\ \penalty 10000\ .$ (7)
In (7), $\chi$ is a Lie generating function acting to normalize the Hamiltonian (6) with respect to all terms of $\mathcal{H}_{1,space}$ not ‘in normal form’ (see subsection 4.2), and $\{\cdot,\cdot\}$ denotes the Poisson bracket operation. The averaging $\cdot$> in the last term of Eq. (7) means to eliminate from $\{\mathcal{H}_{1,space},\chi\}$ all the terms not in normal form.
@@
iii) Our third model considered, $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$, is similar to the secular normal form $\mathcal{H}_{int}^{(1)}$, but obtained within the framework of the octupole approximation of the initial Hamiltonian truncated up to fourth order in the orbital eccentricities. Our interest in such a model, which becomes precise in cases of so-called ‘hierarchical’ systems (planets in nearly circular orbits with $a_{2}/a_{3}<<1$), is motivated by the simplicity in the form of the Hamiltonian, which allows for analytical formulas concerning the dependence of the sequence of bifurcations on the system’s parameters.
@@
Focusing, now, on the above three examples of integrable models approximating the 3D secular dynamics, we implement the geometric method exposed in detail in Section 3 to obtain theoretical predictions for the sequences of bifurcations of periodic orbits arising in each model. New periodic orbits appear stemming from either of the ACR modes A or B, through both saddle-node and pitchfork bifurcations. The sequences are parameterized altering the values of each of the integrals of the energy $\mathcal{E}$ and angular momentum $\sigma_{0}$. The method relies on the use of suitable Hopf variables $\sigma_{1},\sigma_{2},\sigma_{3}$ defined so that, for fixed $\sigma_{0}$, the corresponding reduced manifold in phase space is the 3-sphere $\sigma_{1}^{2}+\sigma_{2}^{2}+\sigma_{3}^{2}=\sigma_{0}^{2}$. All quasi-periodic motions of the system are represented by the curves of transverse intersections of the 2D surfaces of constant energy (e.g. $\mathcal{H}_{int}(\sigma_{0},\sigma_{1},\sigma_{3})=\mathcal{E}$) with a sphere of fixed $\sigma_{0}$. All periodic orbits of the system can then be classified in terms of two different kinds of geometric connections between constant energy surfaces and the spheres of different radii $\sigma_{0}$: i) tangencies, and ii) transverse intersections, where the constant energy condition leads to a surface of degenerate form. Section 3 makes a systematic classification of these cases, the corresponding type of stability of the resulting periodic orbits, as well as the various possibilities for the resulting sequences of bifurcations, depending on the chosen integrable model. Whenever possible, we also provide the corresponding analytical formulas pertinent to the above analysis. Note that our analysis as above is strictly valid in the regular set of initial conditions which avoid the singularities mentioned in [21], [22] (see also [5]). In [21], special motions of the inner body leading to singularities are considered. They could be: i) circular trajectories, ii) coplanar motions and iii) rectilinear motions. In our work, we do not include iii), since it would imply $\mathrm{e}_{2}=1$, while we are interested in motions such that $\mathrm{e}_{2}\in$ $[0,1)$. The other two cases will be discussed in Section 3.2.
@@
To verify the validity of the formulas deduced in section 3, Section 4 implements the method to each of the models $\mathcal{H}_{int}$, $\mathcal{H}_{int}^{(1)}$, or $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$. Besides showing the ability of the method to predict the correct sequence of bifurcations observed in each model, a key result is that even small differences in the chosen model can lead to drastic differences as regards the observed sequence of bifurcations. We explain the geometric origin of such differences. Comparing to the initial non-integrable Hamiltonian (Eq. (3)), we find that the model $\mathcal{H}_{int}^{(1)}$, which represents a higher order normal form for the secular Hamiltonian, yields also the best predictions as regards both the correct thresholds in parameter values and the form of the sequence of bifurcations. Of course, the question of how precisely a secular normal form model can approximate the true dynamics can only be answered through case by case study, i.e, studying various cases of exoplanetary systems. While this is beyond our present scope, we stress that the method exposed in Section 3 is equally applicable to all such cases.
@@ 2.1 Secular Hamiltonian
Following, now, section $2.3$ of [17], we first Jacobi-reduce the Hamiltonian (8), introducing two ‘book-keeping symbols’ ($\varepsilon$, $\eta$) which keeps track of the order of smallness in eccentricity and mutual inclination of all the terms in the Hamiltonian expansion. We briefly recall in the following the main passages of this procedure. We first introduce the following canonical transformation (see also [14])
@@
$R_{2}=L_{2}-H_{2}\,$, $\theta_{r2}=\Omega_{3}-\Omega_{2}\,$,
$R_{3}=L_{2}+L_{3}-H_{2}-H_{3}=\mathrm{AMD}\,$, $\theta_{r3}=-\Omega_{3}\,$,
where $j=2,3$ and
@@
$H_{j}=G_{j}\cos(\mathrm{i}_{j})\,$, $h_{j}=\Omega_{j}\,$,
are the Delaunay variables, and then we express the Hamiltonian given by Eq. (8) as:
$\mathcal{H}_{sec}\equiv\mathcal{H}_{sec}(\Lambda_{2},\Lambda_{3},W_{2},W_{3},R_{2},R_{3},w_{2},w_{3},\theta_{r2}).$
From Hamilton’s equation of the latter, it is easy to recover the constancy of the semi-major axes $a_{2}$, $a_{3}$ and of the $\mathrm{AMD}$. The crucial invariance property of the Hamiltonian is given by the equation $\dot{\theta}_{r2}=(\partial\mathcal{H}_{sec}/\partial R_{2})|_{\theta_{r2}=\pi}=0$, corresponding to the invariance in time of the relation $\theta_{r2}=\Omega_{3}-\Omega_{2}=\pi$ for all trajectories. This means that this substitution can be made directly in the above Hamiltonian. However, after this substitution, while the evolution of the eccentricity vector can be obtained by Hamilton’s equation, the evolution of inclination should be obtained directly from the conservation of the angular momentum (see Eq. (17) of [17]).
In the practical implementation, we can leave the dependence of the Hamiltonian $\mathcal{H}_{sec}$ on the action variables $(\Lambda_{j},W_{j},R_{j})$ through the elements $(a_{j},\mathrm{e}_{j},\mathrm{i}_{j})$ $j=2,3$ and we observe that, after the substitution $\theta_{r2}=\pi$, the Hamiltonian depends on the inclinations only through the trigonometric combination $\cos(\mathrm{i}_{2}+\mathrm{i}_{3})$. This geometrical fact is crucial for the following reduction. In fact, we can introduce the book-keeping control identities:
$\cos(\mathrm{i}_{2})\cos(\mathrm{i}_{3})\!=\!\varepsilon^{2}\eta\left(\cos(\mathrm{i}_{2})\cos(\mathrm{i}_{3})-1\right)+1\,$, $\sin(\mathrm{i}_{2})\sin(\mathrm{i}_{3})\!=\!\varepsilon^{2}\eta\sin(\mathrm{i}_{2})\sin(\mathrm{i}_{3})\,$, (12)
where $\mathcal{O}(\varepsilon^{2})$ stands for ‘second order in the eccentricities and inclinations’ while $\eta$ is the book-keeping symbol separating the Hamiltonian terms depending on powers of the quantity $\cos(\mathrm{i}_{2}+\mathrm{i}_{3})$ from those not depending on the mutual inclination $\mathrm{i}_{mut}$. Moreover, we use the substitution rule $\sin(\mathrm{i}_{2})\sin(\mathrm{i}_{3})=\cos(\mathrm{i}_{2})\cos(\mathrm{i}_{3})+F(a_{2},a_{3},\mathrm{e}_{2},\mathrm{e}_{3};L_{z})$ (where the function $F$ is explicitly reported in Eq. (19) of [17]), determined by the constancy of the angular momentum. Substituting the above expressions into the Hamiltonian, and truncating the resulting expression up to a preselected maximum order in book-keeping $N_{bk}$, by symmetry the terms in equal powers of the products $\sin(\mathrm{i}_{2})\sin(\mathrm{i}_{3})$ and $\cos(\mathrm{i}_{2})\cos(\mathrm{i}_{3})$ are opposite, and thus they are canceled. Hence, the Hamiltonian resumes the form:
$\mathcal{H}_{sec}=\sum_{s_{1}=0}^{N_{bk}/2}\eta^{s_{1}}\varepsilon^{2s_{1}}\mathcal{H}_{sec,s_{1}}(\mathrm{e}_{2},\mathrm{e}_{3},w_{2},w_{3};a_{2},a_{3},L_{z})\penalty 10000\ \penalty 10000\ .$ (13)
We finally set $\varepsilon=1$ and $\eta=1$ (as implied by the identities (12)) of the book-keeping parameters, and write the truncated (up to book-keeping order $N_{bk}$) Hamiltonian as:
$\mathcal{H}_{sec}=\mathcal{H}_{planar}+\mathcal{H}_{space}$ (14)
where
$\mathcal{H}_{planar}=\mathcal{H}_{sec,0}(\mathrm{e}_{2},\mathrm{e}_{3},w_{2}-w_{3};a_{2},a_{3})\,$$\qquad$ $\mathcal{H}_{space}=\sum_{s_{1}=1}^{N_{bk}/2}\mathcal{H}_{sec,s_{1}}(\mathrm{e}_{2},\mathrm{e}_{3},w_{2},w_{3};a_{2},a_{3},L_{z})\penalty 10000\ .$ (15)
Finally, we expand the Hamiltonian in the orbital eccentricities, as explained in Step $3$ of [17]. The angle $\theta_{r3}$ is ignorable, hence, the action variable $R_{3}$ is constant of motion. Considering $R_{3}$ as a parameter, we can substitute it in terms of the (also constants) $(L_{2},L_{3},L_{z})$ via the relation $R_{3}=L_{2}+L_{3}-L_{z}$. Then, we obtain the following decomposition of the Hamiltonian:
$\mathcal{H}_{sec}(w_{2},w_{3},W_{2},\,W_{3};L_{z})=\mathcal{H}_{planar}(w_{2}-w_{3},W_{2},W_{3})+\mathcal{H}_{space}(w_{2},w_{3},W_{2},W_{3};L_{z})\,.$ (16)
The Hamiltonian can now be formally considered as with 2 DOF. Furthermore, the spatial part $\mathcal{H}_{space}$ can be split as $\mathcal{H}_{space}=\mathcal{H}_{0,{space}}+\mathcal{H}_{1,{space}}$, where $\mathcal{H}_{0,{space}}$ contains terms independent of the angles or depending trigonometrically on the difference $w_{2}-w_{3}=\omega_{3}-\omega_{2}$, and $\mathcal{H}_{1,{space}}$ contains terms depending trigonometrically on combinations of the angles $w_{2},w_{3}$ other than their difference. Then we write
@@
The first two terms in the above expression give rise to the Hamiltonian
@@
where the $\mathrm{AMD}$ is defined as in Eq. (2). The Hamiltonian $\mathcal{H}_{int}$ admits a second integral of motion $J=(W_{2}+W_{3})/2$. After a linear symplectic change of variables (see Eq. (26) in subsection 3.1 below), the Hamiltonian can be formally regarded as 1 DOF. The fixed points of the 1 DOF Hamiltonian actually correspond to periodic orbits of the original symplectic flow of $\mathcal{H}_{int}$ in all four variables $(w_{2},w_{3},W_{2},W_{3})$. The bifurcation chains and stability of such periodic orbits are our main subject of interest in the sequel.
@@ 2.2 Numerical phase portraits
$\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})=\bigg\{$ $(X_{2},Y_{2},X_{3},Y_{3})\in\mathbb{R}^{4}:\penalty 10000\ \mathcal{H}_{sec}(X_{2},Y_{2},X_{3},Y_{3}=0;\mathrm{AMD})=\mathcal{E}\,,\,Y_{3}=0\,$, (19)
@@
$X_{j}=-\sqrt{2\,W_{j}}\cos(w_{j})\,$, $Y_{j}=\sqrt{2\,W_{j}}\sin(w_{j})\,$, $j=2,3\penalty 10000\ \penalty 10000\ .$ (20)
@@
$\cos(\mathrm{i}_{mut})$ $=\frac{L_{z}^{2}-\Lambda_{2}^{2}-\Lambda_{3}^{2}+\Lambda_{2}^{2}\,\mathrm{e}_{2}^{2}+\Lambda_{3}^{2}\,\mathrm{e}_{3}^{2}}{2\Lambda_{2}\Lambda_{3}\sqrt{1-\mathrm{e}_{2}^{2}}\sqrt{1-\mathrm{e}_{3}^{2}}}=\frac{L_{z}^{2}-G_{2}^{2}-G_{3}^{2}}{2G_{2}G_{3}}\penalty 10000\ \penalty 10000\$(21)
@@
$\mathrm{i}_{max}=\cos^{-1}\left(\frac{L_{z}^{2}-\Lambda_{2}^{2}-\Lambda_{3}^{2}}{2\Lambda_{2}\Lambda_{3}}\right)\penalty 10000\ \penalty 10000\ .$ (22)
@@
Figure 1: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ in the plane $(X_{2},Y_{2})$ with $\mathrm{AMD}$ fixed and different values of energy. The surfaces of section are computed by a numerical integration of trajectories in the Hamiltonian truncated at multipolar degree $N_{P}=5$, order $N_{bk}=12$ in the eccentricities, and energies (from left to right) $\mathcal{E}=-6.77\cdot 10^{-5},-2.53\cdot 10^{-5},-1.92\cdot 10^{-5},{-1.73\cdot 10^{-5},}-1.17\cdot 10^{-5}$$2.61\cdot 10^{-6}$.
@@
Figure 1 allows to recognize one out of all in-principle possible chains of bifurcations that can emerge in 3D secular models of the type discussed above. Such models may differ in the parameters (masses and semi-major axes of the planets, value of the AMD) as well as on the truncation orders (in the multipoles and/or the orbital eccentricities and inclinations). Our aim in the next section is to provide a systematic method to predict which out of all possibilities actually takes place in a system with given parameters. In the case of Fig. 1, a quick description of the numerically obtained sequence of bifurcations is as follows: for low $\mathcal{E}$, i.e., for nearly planar orbits (left panel in the first row of Fig. 1), the phase portrait exhibits two stable fixed points which correspond to the ACR modes A (right fixed point) and B (left fixed point, reversed with respect to the left-right position of the same fixed points in figure 9 of [17]). In between the centers A and B we have a continuous transition of quasi-periodic orbits represented in the surface of section by closed curves surrounding either the fixed point A or B. No unstable points or separatrices exist between the stable points, while the apparent separation of the two domains is only an artifact of the projection of the phase space, whose real topology is the one of the 2-sphere (see below), to the plane $(X_{2},Y_{2})$. The second panel in the first row of Fig. 1 shows, now, the birth of the families called ‘C1’ (unstable) and ‘C2’ (stable) in [17] through a saddle-node bifurcation. We denote by $\mathcal{E}_{C}$ the critical value of the energy in which the saddle-node bifurcation takes place. Increasing the value of the energy $\mathcal{E}$, we see (third panel of first row) that the unstable orbit C1 becomes stable through a pitchfork bifurcation giving rise to two ‘off-axis’ ($Y_{2}\neq 0$) unstable orbits , namely $D_{1}$ and $D_{2}$, which (as the energy $\mathcal{E}$ further increases) join each other through an inverse pitchfork bifurcation turning the left fixed point of the middle panel of Fig. 1 from stable to unstable. More precisely, at a critical value of the energy, the fixed points $D_{1}$ and $D_{2}$ collide with the B-mode, who becomes unstable (see first panel, second row). For still higher energy $\mathcal{E}$, the C2 orbit finally turns also to unstable via the ‘Lidov-Kozai’ mechanism. We denote by $\mathcal{E}_{C,2}$ the critical value of the energy leading to the Lidov-Kozai instability. In the Lidov-Kozai regime, a large volume of the phase space is dominated by chaotic motions, excluded by any of the integrable models used in the sequel to approximate the flow under the complete Hamiltonian. We call transition regime the one holding at energies in the interval $\mathcal{E}_{C}\leq\mathcal{E}\leq\mathcal{E}_{C,2}$, representing the transition between the planar-like regime and the Lidov-Kozai one (see section 3.3 of [17]).